solve the heat conduction equation: 8*∂^2u(x,t) /∂x^2 = ∂u(x,t)/∂t; for0<x<5 and t>0 the following boundary and initial conditions: u(0,t) = u(5,t) = 0, u(x,0) = 2sin(πx) − 4sin(2πx)
1
Expert's answer
2019-03-27T11:30:14-0400

The solution of this equation will be sought in the form


"u(x,t)=X(x)T(t)\rightarrow\frac{\partial u}{\partial t}=\frac{\partial}{\partial t}\left(X(x)T(t)\right)=X(x)T'(t)""\frac{\partial^2 u}{\partial x^2}=\frac{\partial^2}{\partial x^2}\left(X(x)T(t)\right)=X''(x)T(t)"

Substitute the derived derivatives in the initial equation


"8\cdot\frac{\partial^2 u}{\partial x^2}=\frac{\partial u}{\partial t}\rightarrow\left.X(x)T'(t)=8\cdot X''(x)T(t)\right|\div\left(8\cdot X(x)T(t)\right)\rightarrow""\frac{T'(t)}{8\cdot T(t)}=\frac{X''(x)}{X(x)}\rightarrow\boxed{\frac{X''(x)}{X(x)}=-\lambda=\frac{T'(t)}{8\cdot T(t)}}"

Rewrite the specified boundary conditions through the entered functions X(x) and T(t)


"u(0,t)=X(0)T(t)=0\rightarrow X(0)=0\\
u(5,t)=X(5)T(t)=0\rightarrow X(5)=0"

Conclusion


"\left\lbrace\begin{array}{c}
8\cdot u_{xx}=u_t\\
u(x,0)=2\sin(\pi x)-4\sin(2\pi x)\\
u(0,t)=u(5,t)=0\end{array}\right.\rightarrow""\left\lbrace\begin{array}{c}
\displaystyle\frac{X''(x)}{X(x)}=-\lambda=\displaystyle\frac{T'(t)}{8\cdot T(t)}\\[0.5 cm]
u(x,0)=2\sin(\pi x)-4\sin(2\pi x)\\
X(0)=0\\
X(5)=0\end{array}\right."

To find the eigenvalues and functions, it is necessary to solve the Sturm-Liouville problem


"\boxed{\left\lbrace\begin{array}{c}
X''+\lambda X=0\\
X(0)=0\\
X(5)=0\end{array}\right.}"

We can immediately assume that it only


"\lambda=n^2>0"

gives a non-trivial solution.

The solution is sought in the form


"X(x)=e^{kx}\rightarrow X''=k^2\cdot e^{kx}"

Then,


"X''+\lambda X=0\rightarrow k^2\cdot e^{kx}+n^2\cdot e^{kx}=0\rightarrow e^{kx}\cdot\left(k^2+n^2\right)=0""k^2=-n^2\rightarrow
\left\lbrack\begin{array}{c}
k_1=\sqrt{-n^2}=in\\
k_2=-\sqrt{-n^2}=-in\\
\end{array}\right."

Conclusion,


"X(x)=C_1\cdot e^{k_1x}+C_2\cdot e^{k_2x}=C_1\cdot e^{inx}+C_2\cdot e^{-inx}\rightarrow"




"X(x)=A_1\cdot\cos(nx)+A_2\cdot \sin(nx)"

"\boxed{X(x)= A_1\cdot\cos(nx)+A_2\cdot \sin(nx)}"

To find constants A_1 and A_2 we use the boundary conditions.


"X(0)=0=A_1\cos(n\cdot 0)+A_2\sin(n\cdot 0)\rightarrow\boxed{A_1=0}""X(5)=0=A_2\cdot\sin(5n)"

Since we are trying to find a non-trivial solution, then


"\sin(5n)=0\rightarrow 5n=\pi k,\quad k\in\mathbb{N}\rightarrow\boxed{n_k=\left(\frac{\pi k}{5}\right),\quad k\in\mathbb{N}}"

Conclusion,


"\boxed{\left\lbrace\begin{array}{c}
\lambda_k=\displaystyle\left(\frac{\pi k}{5}\right)^2\\
X_k(x)=A_2\cdot\sin\left(\displaystyle\frac{\pi kx}{5}\right)
\end{array}\right.}"

Then,


"\frac{T'(t)}{8\cdot T(t)}=-\lambda_k\rightarrow\frac{dT}{T}=-8\cdot\lambda_kdt\rightarrow""\int\frac{dT}{t}=\int(-8\cdot\lambda_kdt)\rightarrow\ln|T_k(t)|=\ln|C_k|-8\lambda_kt\rightarrow""\boxed{T_k(t)=C_k\cdot e^{-8\cdot\lambda_kt}}"

General conclusion,


"\boxed{u(x,t)=\sum_{k=1}^{\infty}\left(B_k\cdot e^{-8\cdot\lambda_kt}\cdot\sin\left(\frac{\pi kx}{5}\right)\right)}"

To find constants B_k we use the boundary conditions.


"u(x,0)=2\sin(\pi x)-4\sin(2\pi x)=\rightarrow""\sum_{k=1}^{\infty}\left(B_k\cdot\sin\left(\frac{\pi kx}{5}\right)\right)=2\sin(\pi x)-4\sin(2\pi x)\rightarrow""\boxed{\left\lbrace\begin{array}{c}
k=5 : B_5=2\\
k=10 : B_{10}=-4\\
B_k\equiv 0,\quad k=\not 5,10
\end{array}\right.}"

General conclusion,


"u(x,t)=B_5\cdot e^{-8\cdot\lambda_5 t}\cdot\sin(\pi x)+B_{10}\cdot e^{-8\cdot\lambda_{10} t}\cdot\sin(2\pi x)""=2\cdot e^{-8\pi^2\cdot 5^2\cdot t/25}\cdot\sin(\pi x)-4\cdot e^{-8\pi^2\cdot 10^2\cdot t/25}\cdot\sin(2\pi x)\rightarrow""\boxed{u(x,t)=2\cdot e^{-8\pi^2 t}\cdot\sin(\pi x)-4\cdot e^{-32\pi^2t}\cdot\sin(2\pi x)}"

ANSWER

"u(x,t)=2\cdot e^{-8\pi^2 t}\cdot\sin(\pi x)-4\cdot e^{-32\pi^2t}\cdot\sin(2\pi x)"